OpenStudy (anonymous):

derivate y=x(2x+1)^3

OpenStudy (amistre64):

looks like i see a product, a power, and a chain rule ....

OpenStudy (anonymous):

yes, I used the chain rule

OpenStudy (amistre64):

did you use the other rules as well? lets see what youve done

OpenStudy (anonymous):

I have this( X). 6(2x+1)+ (2x+1)^3.(1)

OpenStudy (amistre64):

y=x(2x+1)^3 D[y] = D[x (2x+1)^3] product D[y] = D[x] (2x+1)^3 + x D[(2x+1)^3] power/chain D[y] = D[x] (2x+1)^3 + x 3(2x+1)^2 D[2x+1]

OpenStudy (anonymous):

The answer is DY= (2x+1)^2 (8x+1) But i dont know (8x+1)

OpenStudy (amistre64):

you understand how i got this far? D[y] = D[x] (2x+1)^3 + x 3(2x+1)^2 D[2x+1]

OpenStudy (amistre64):

D[y] = D[x] (2x+1)^3 + x 3(2x+1)^2 D[2x+1] D[y] = y' D[x] = 1 y' = (2x+1)^3 + x 3(2x+1)^2 D[2x+1] ^^^^^^ just a simple derivative y' = (2x+1)^3 + x 3(2x+1)^2 (2) y' = (2x+1)^3 + 6x(2x+1)^2 we can factor out (2x+1)^2 ... this is just algebra y' = (2x+1)^2 ( 2x +1 + 6x) etc ...

OpenStudy (anonymous):

Thanks

OpenStudy (amistre64):

yep